Find the Arihmetic means of the sequence
23, ___, ____, ____, 67

Answer by stanbon(75887)

Find the Arihmetic means of the sequence
23, ___, ____, ____, 67
(23+67)/2=45
(23+45)/2=34
(45+67)/2=56
Sequence: 23,34,45,56,67
Cheers,
Stan H.